OpenStudy (anonymous):

Choose the correct product of (5x − 9)2

OpenStudy (anonymous):

??

OpenStudy (anonymous):

(5x-9)^2=(5x-9)(5x-9)=(25x^2-45x-45x+81)=25^x2-90x+81!..... Happy Studying!!!!!

OpenStudy (anonymous):

actually...../

OpenStudy (anonymous):

25x^2-90x+81...@typo